OpenStudy (anonymous):

How do I find the distance from (6,1) to the line defined by y=3x-7. Express as a radical or a number rounded to the nearest hundredth.

OpenStudy (anonymous):

This is what I would do, but there might be an easier way: find the line perpendicular to y=3x-7 that goes through (6,1). Than find the intersection, (a,b), of those two lines. The distance then will be sqrt((6-a)^2+(1-b)^2)

OpenStudy (angela210793):

\[d=|\sqrt{Ax1+By1+C}|/\sqrt{A^2+B^2}\]

OpenStudy (anonymous):

Now the shortest distance is when Lines are parallel. y = 3x - C, passes through (6,1) 1 = 18 -C C=17 Equation becomes y = 3x -17 Now there is a formula to find distance between parallel lines. \[d = \frac{|C_1 -C_2|}{\sqrt{a^2 +b^2}}\] ax + by + c = 0 is the standard form. line equation is standard form y -3x +7 = 0 and y -3x -17 = 0 a = -3 b = 1 \[C_1 = 7\] \[C_2 = -17\] \[d = \frac{17 + 7 }{\sqrt{9 +1}}\] \[d = \frac{24}{\sqrt{10}}\]